Dr. David Bryan Carpenter

Northeast Parallel Architectures Center 111 College Place Syracuse University Syracuse, NY 13244-4100 Phone: (315) 443-5068 FAX : (315) 443-1973 email: dbc@npac.syr.edu

March 14, 2000

Recommendation for Sang Boem Lim

Dear Colleague,

I have supervised Sang Boem Lim in activities related to our research group since June 1998, when he joined us as an “independent study” student. Sang Lim demonstrated great promise, and in January 1999 he became a Ph.D. student, and a Graduate Resarch Assistant in my group.

Sang Lim has been essentially involved with the implementation of our mpiJava system—a Java binding of MPI developed as part of the ongoing HPJava project. He was particularly involved in the work to introduce object serialization into the mpiJava API, and developed software that was the main basis for several recent workshop and conference papers, on which he was a co-author (“Automatic Object Serialization in the mpiJava Interface to MPI” in proceedings 3rd MPI Developers’ and Users’ Conference MPI Software Technology Press 1999; “mpiJava: An Objectoriented Java Interface to MPI”, in proceedings Intl. Workshop on Java for Parallel and Distributed Computing, IPPS/SPDP ’99; “Object Serialization for Marshalling Data in a Java Interface to MPI”, in proceedings ACM 1999 Java Grande Conference, ACM Press 1999).

The mpiJava software for which Sang Lim is now primarily responsible for has been available on the Web for over a year. It is regularly downloaded by people wanting to use it for research and teaching (about 30 downloads per month, currently). The software is a cornerstone of an NSF-ITR proposal recently submitted on behalf of FSU (“ITR/ACS:Integration of High Performance Message-based Programming Environments with Distributed Objects and the Web”). It has also served as a reference implementation in a recent Java message-passing standardization effort, attempted by a working group of the Java Grande Forum.

Besides his work on the underlying system software, Sang Lim has produced mpiJava demos for various conferences. Currently he is studying the Jini and Javaspaces software from Sun, which we expect will be important to future work on scalable fault-tolerant computing, following on from the mpiJava work. Sang Lim is an exceptional student; he will be a vital resource for continuing work on Java-based parallel computing at FSU.

Yours sincerely,

Bryan Carpenter